5.  Neptune Mining (NM) has two outstanding bonds:

Bond 1:  Par Value = $200 million; Coupon Rate = 3.1%; Yield to Maturity = 4.2%; Maturity Date = 40 years

Bond 2:  Par Value = $300 million; Coupon Rate during Years 1-8 = 4.3%; Coupon Rate during Years 9-10 = 4.8%; Yield to Maturity = 4.1%

Details on NM equity shares:

Current Dividend = $1.58; Growth Rate in Dividends = 2.83%; Required return = 9.61%

A.  Compute the price of both bonds.

B.  Compute the price of the stock using the Gordon growth model.

C.  Compute the sinking fund payment for Bond 1 if the return provided by the financial institution is 4.2% and the payments begin exactly 15 years before the bond matures.
Answer #1

A. Prices of bonds.
Bond 1:

Bond 1 Calculation
Future Value 200000000 Given
Coupon 6200000 3.1%*200000000
YTM 4.20% Given
Time to maturity 40 Given
Periodicity Annual Given
Price 15,77,22,379.80 Financial Calculator


N = 40
PMT = 6200000
FV = 200000000
I/Y = 4.2
CPT PV = 157722379.8

Excel can also be used:

Year Cash Flow Discounting Factor Present Value
0 0 1 0
1 6200000 0.959692898 5950095.97
2 6200000 0.921010459 5710264.85
3 6200000 0.883887197 5480100.62
4 6200000 0.848260266 5259213.65
5 6200000 0.814069353 5047229.99
6 6200000 0.781256577 4843790.77
7 6200000 0.749766388 4648551.61
8 6200000 0.719545478 4461181.96
9 6200000 0.690542685 4281364.65
10 6200000 0.662708911 4108795.25
11 6200000 0.635997036 3943181.62
12 6200000 0.610361838 3784243.4
13 6200000 0.585759922 3631711.51
14 6200000 0.562149637 3485327.75
15 6200000 0.539491014 3344844.29
16 6200000 0.517745695 3210023.31
17 6200000 0.496876867 3080636.57
18 6200000 0.4768492 2956465.04
19 6200000 0.457628791 2837298.5
20 6200000 0.439183101 2722935.23
21 6200000 0.421480903 2613181.6
22 6200000 0.404492229 2507851.82
23 6200000 0.38818832 2406767.58
24 6200000 0.372541574 2309757.76
25 6200000 0.357525503 2216658.12
26 6200000 0.343114686 2127311.05
27 6200000 0.329284727 2041565.31
28 6200000 0.316012214 1959275.73
29 6200000 0.303274678 1880303
30 6200000 0.291050555 1804513.44
31 6200000 0.27931915 1731778.73
32 6200000 0.268060605 1661975.75
33 6200000 0.257255859 1594986.32
34 6200000 0.246886621 1530697.05
35 6200000 0.236935337 1468999.09
36 6200000 0.22738516 1409787.99
37 6200000 0.218219923 1352963.52
38 6200000 0.20942411 1298429.48
39 6200000 0.200982832 1246093.56
40 206200000 0.192881796 39772226.4
Bond Price 157722380

Bond 2:

Bond 1 Calculation
Future Value 30,00,00,000.00 Given
Coupon 1-8      1,29,00,000.00 4.3%*300000000
Coupon 8-10      1,44,00,000.00 4.8%*300000000
YTM 4.10% Given
Time to maturity 10 Given
Periodicity Annual Given
Price 10,61,60,084.62 Financial Calculator

Excel:

Year Cash Flow Discounting Factor Present Value
0 0 1 0
1      1,29,00,000.00 0.960614793 12391930.8
2      1,29,00,000.00 0.922780781 11903872.1
3      1,29,00,000.00 0.88643687 11435035.6
4      1,29,00,000.00 0.851524371 10984664.4
5      1,29,00,000.00 0.817986907 10552031.1
6      1,29,00,000.00 0.785770324 10136437.2
7      1,29,00,000.00 0.754822598 9737211.51
8      1,29,00,000.00 0.725093754 9353709.42
9      1,44,00,000.00 0.696535786 10030115.3
10      1,44,00,000.00 0.669102581 9635077.16
Price 106160085

B. Price of stock

Dividend 1.58
Growth rate 2.83%
Rate of return 9.61%
Stock price 23.96333333

Formula:
[ D0*(1+g) ] / (r - g)

where D0 is the current dividend
g = growth rate
r = rate of return

C. Annual payment to Sinking Fund:

Redemption 206200000
Time to Maturity 15
Rate 4.20%
Annual Sinking Fund payment (1,01,45,747.70)

PMT function of excel was used.
=PMT(4.2%,15,,206200000)

Using a financial calculator:

I/Y = 4.2%
N = 15
FV = 206200000
CPT PMT = 10145747.7